OpenStudy (anonymous):

-4 3/5x 3 1/6. explain? would it be -12 3/30?

OpenStudy (anonymous):

\[-4\frac{ 3 }{ 5 }*3\frac{ 1 }{ 6 }=-1\frac{ 23 }{ 30 }\]

OpenStudy (anonymous):

lol thank you